OpenStudy (anonymous):

There are 6 finalists in a writing competition. Of the finalists, 2 will be awarded trophies. How many ways are there for the 2 trophy winners to be chosen?

OpenStudy (anonymous):

6 * 5 = 30 6 finalists for the 1st trophy and then 5 finalists for the 2nd trophy.

OpenStudy (anonymous):

6 P 2 = 6!/(6-2)!= 720/24 = 30